OpenStudy (anonymous):

can anyone help me with this problem 8x5-15x3

OpenStudy (anonymous):

8x5-15x3 40-45 -5

OpenStudy (anonymous):

\[8x ^{5}-15x ^{3}\]

OpenStudy (anonymous):

(8x2-15)(x3) (8x2-15)(x3)=0 x3=0.........so x=0 8x2-15=0 8x2=15 divide by 8 both sides x2=15/8 x=sqrt of 15/8